OpenStudy (anonymous):

Ok here is my main question, The answer to 4 3/5 plus 3 4/5 is= 7 7/5. So you know how that is a inproper fraction, right? Well how do you make the denaminator bigger than the numinator?

OpenStudy (nowhereman):

You should observe how such an improper fraction is formed and what its meaning a b/c actually means a + b/c so you can write 7/5 as (1*5+2)/5

OpenStudy (anonymous):

Why * (Multiply) I tought you said a b/c means a+(plus) b/c

OpenStudy (nowhereman):

yes, but observe the parentheses, you still have to rewrite this a bit

OpenStudy (anonymous):

ok thanks but how do you reduce 7 7/5?

OpenStudy (nowhereman):

so, did you get something for 7/5?

OpenStudy (anonymous):

Uhm... How do you reduce the inproper fraction? to a fraction a normal fraction?

OpenStudy (anonymous):

Btw the answer to your question is 49/5

OpenStudy (nowhereman):

if you want to convert an improper to a normal fraction a b/c you have to write the integral part as (a*c)/c so you get a b/c = (a*c+b)/c

OpenStudy (anonymous):

hello ?

OpenStudy (nowhereman):

sorry?

OpenStudy (anonymous):

I still dont get it =/! My dad and I are trying to figure my homework out....

OpenStudy (anonymous):

Can you do a number example please?

OpenStudy (nowhereman):

So if you have 5 2/3 you get \[5 = 5\cdot \frac{3}{3} = \frac{5\cdot 3}{3} = \frac{15}{3}\] and together it is \[5\frac{2}{3} = 5 + \frac{2}{3} = \frac{15}{3} + \frac 2 3 = \frac{15+2}{3} = \frac{17}{3}\]